OpenStudy (anonymous):

Solve by completing the square: x2 + 8x = -3.

hero (hero):

since b = 8, then \((b/2)^2 = (8/2)^2 = (4)^2 = 16\) So add 16 to both sides to get x^2 + 8x + 16 = 16 - 3 (x + 8)^2 = 13 Finish solving for x.